I am trying to understand how x - .30x turns into .7
This is only a part of the question I am working on which is 130.90=x - .30x
I am stumped on the x - .30x part ...this question is an example question and provides the ultimate answer but does not explain
I hope you can help me
Thanks

Answer by Alan3354(69443)

I am trying to understand how x - .30x turns into .7
It doesn't.
x - 0.30x = 0.7x, not 0.7
---
x - 0.3x = 1x - 0.3x = x*(1 - 0.3) = 0.7x
--------
This is only a part of the question I am working on which is 130.90=x - .30x
I am stumped on the x - .30x part ...this question is an example question and provides the ultimate answer but does not explain
-------------
130.90=x - .30x
130.90 = 0.7x
x = 130.9/0.7
x = 187